The log Henry's law constant (units of L-atm/mole
Answer #1

Solution:

The dimenssionless henrys law constant is given as,

KH(water/Air) = KH/RT   

Where, KH(water/Air) - Dimenssionless henrys law constant

KH - Logarithmic henrys law constant L - atm/mole

R - ideal gas constant

T - temperature

As it is mentioned in the problem log henrys law constant is given, to convert it into constants we have to take 10 to the power so that it will get as,

logKH = 10LogKH

Now find out values for tricloroethylene = 101.03 = 10.71 l-atm/mol

for tetrachloroethylene = 101.44 = 27.54

for 1,2 - dimethylbenzene - 100.71 = 5.12

for parathion = 10-3.42 = 3.8 * 10-4

Now put these values one by one in equation 1 to calculate KH dimenssionless value,

For trichloroethylene:

KH(trichloroethylene) = 10.71/(0.08205*298)

R = 0.08206 L-atm/mole

T = 25+273 = 298 K

KH(trichloroethylene) = 0.438

For Tetrachloroethylene

KH(tetrachloroethylene) = 27.54/(0.08206 * 298) = 1.1

For 1,2 dimethylbenzene

KH(1,2 - dimethylbenzene) = 5.12/(0.08206*298) = 0.21

For Parathion

KH(parathion) = 3.8 * 10-4/(0.08206 * 298) = 1.6*10-5

Now to decide ease of stripping from water to airon the basis of henrys law constant.

Henrys law constant increases ease of stripping decrreases means it is harder to strip water into air.

therefore,

Parathion>1,2-dimethylbenzene>trichloroethylene>tetrachloroethylene